# Online Posterior Sampling with a Diffusion Prior

### v1

@@ Abstract
Posterior sampling in contextual bandits with a Gaussian prior can be implemented exactly or approximately using the Laplace approximation. The Gaussian prior is computationally efficient but it cannot describe complex distributions. In this work, we propose approximate posterior sampling algorithms for contextual bandits with a diffusion model prior. The key idea is to sample from a chain of approximate conditional posteriors, one for each stage of the reverse process, which are estimated in a closed form using the Laplace approximation. Our approximations are motivated by posterior sampling with a Gaussian prior, and inherit its simplicity and efficiency. They are asymptotically consistent and perform well empirically on a variety of contextual bandit problems.
@@ 1 Introduction
A multi-armed bandit [26, 6, 29] is an online learning problem where an agent sequentially interacts with an environment over $n$ rounds with the goal of maximizing its rewards. In each round, it takes an action and receives its stochastic reward. The mean rewards of the actions are unknown a priori and must be learned. This leads to the exploration-exploitation dilemma: explore actions to learn about them or exploit the action with the highest estimated reward. Bandits have been successfully applied to problems where uncertainty modeling and adaptation to it are beneficial, such as in recommender systems [31, 53, 24, 34] and hyper-parameter optimization [33].
@@
Contextual bandits [28, 31] with linear [13, 1] and generalized linear models (GLMs) [16, 32, 2, 25] have become popular due to the their flexibility and efficiency. The features in these models can be hand-crafted or learned from historic data [39], and the models can be also updated incrementally [1, 23]. While the original algorithms for linear and GLM bandits were based on upper confidence bounds (UCBs) [13, 1, 16], Thompson sampling (TS) is more popular in practice [11, 3, 41, 43]. The key idea in TS is to explore by sampling from the posterior distribution of model parameter $\theta_{*}$. TS uses the prior knowledge about $\theta_{*}$ to speed up exploration [11, 39, 35, 9, 20, 19, 5]. When the prior is a multivariate Gaussian, the posterior of $\theta_{*}$ can be updated and sampled from efficiently [11]. Unfortunately, this prior has limited expressive power, because it cannot even represent multimodal distributions. To address this limitation, we study posterior sampling with complex priors.
@@
We make the following contributions. First, we propose novel posterior sampling approximations for linear models and GLMs with a diffusion model prior. The key idea is to sample from a chain of approximate conditional posteriors, one for each stage of the reverse process, which are estimated in a closed form. In linear models, each conditional is a product of two Gaussians, representing prior knowledge and diffused evidence (Theorem 2). In GLMs, each conditional is obtained by a Laplace approximation, which mixes prior knowledge and evidence (Theorem 4). Our approximations are motivated by posterior sampling with a Gaussian prior, and inherit its simplicity and efficiency. The prior works (Section 7) rely on the score of the likelihood of observations. In online learning, the likelihood score tends to infinity as the number of observations increases, which causes divergence of the existing approximations (Section 6.2).
@@
Our second major contribution is in theory. More specifically, we prove that our posterior approximation is asymptotically consistent (Theorem 3). The key idea in the proof is that the conditional posteriors in Theorem 2 concentrate at a scaled unknown model parameter as the number of observations increases. While this claim is asymptotic, it is an expected property of a posterior distribution. Many prior works, such as Chung et al. [12], do not poses this property.
@@
Our last contribution is empirical evaluation on contextual bandit problems. We focus on bandits because the ability to represent all levels of uncertainty well is critical for exploration. Our experiments (Section 6.2) show that a score-based method fails to do so. Note that our posterior approximations are general and not restricted to bandits.
@@ 2 Setting
We start with introducing our notation. Random variables are capitalized, except for Greek letters like $\theta$. We denote the marginal and conditional probabilities under probability measure $p$ by $p(X=x)$ and $p(X=x\mid Y=y)$, respectively. When the random variables are clear from context, we write $p(x)$ and $p(x\mid y)$. We denote by $X_{n:m}$ and $x_{n:m}$ a collection of random variables and their values, respectively. For a positive integer $n$, we define $[n]=\left\{1,\dots,n\right\}$. The indicator function is $\mathds{1}\!\left\{\cdot\right\}$. The $i$-th entry of vector $v$ is $v_{i}$. If the vector is already indexed, such as $v_{j}$, we write $v_{j,i}$. We denote the maximum and minimum eigenvalues of matrix $M\in\mathbb{R}^{d\times d}$ by $\lambda_{1}(M)$ and $\lambda_{d}(M)$, respectively.
@@
where $g:\mathbb{R}\to\mathbb{R}$ is the mean function and $\varepsilon_{\ell}$ is independent zero-mean $\sigma^{2}$-sub-Gaussian noise for $\sigma>0$. Let $p(h\mid\theta_{*})$ be the likelihood of observations in history $h$ under model parameter $\theta_{*}$ and $p(\theta_{*})$ be its prior probability. By Bayes’ rule, the posterior distribution of $\theta_{*}$ given $h$ is
@@ 2.1 Linear Model
The posterior of $\theta_{*}$ in linear models has the following form. Let $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$ be the prior distribution, where $\theta_{0}\in\mathbb{R}^{d}$ and $\Sigma_{0}\in\mathbb{R}^{d\times d}$ are the prior mean and covariance, respectively.
@@
Let $g$ in (1) be an identity and $\varepsilon_{\ell}\sim\mathcal{N}(0,\sigma^{2})$. Then the likelihood of $h$ under model parameter $\theta_{*}$ is $p(h\mid\theta_{*})\propto\exp[-\sigma^{-2}\sum_{\ell=1}^{N}(y_{\ell}-\phi_{\ell}^{\top}\theta_{*})^{2}]$.
Under these assumptions, $p(\theta_{*}\mid h)\propto\mathcal{N}(\theta_{*};\hat{\theta},\hat{\Sigma})$ [10], where
@@
$\bar{\theta}=\sigma^{-2}\bar{\Sigma}\sum_{\ell=1}^{N}\phi_{\ell}y_{\ell}$, and $\bar{\Sigma}^{-1}=\sigma^{-2}\sum_{\ell=1}^{N}\phi_{\ell}\phi_{\ell}^{\top}$. Therefore, the posterior of $\theta_{*}$ is a product of two multivariate Gaussians, representing the prior knowledge about $\theta_{*}$, given by $\mathcal{N}(\theta_{0},\Sigma_{0})$; and empirical evidence, given by $\mathcal{N}(\bar{\theta},\bar{\Sigma})$.
@@ 2.2 Generalized Linear Model
Algorithm 1 $\tt IRLS$: Iteratively reweighted least squares.
@@
5:     for stage $\ell=1,\dots,N$ do $z_{\ell}\leftarrow\phi_{\ell}^{\top}\hat{\theta}+(y_{\ell}-g(\phi_{\ell}^{\top}\hat{\theta}))/\dot{g}(\phi_{\ell}^{\top}\hat{\theta})$
6:     $\hat{\Sigma}\leftarrow\left(\Sigma_{0}^{-1}+\sum_{\ell=1}^{N}\dot{g}(\phi_{\ell}^{\top}\hat{\theta})\phi_{\ell}\phi_{\ell}^{\top}\right)^{-1}\,,\ \hat{\theta}\leftarrow\hat{\Sigma}\left(\Sigma_{0}^{-1}\theta_{0}+\sum_{\ell=1}^{N}\dot{g}(\phi_{\ell}^{\top}\hat{\theta})\phi_{\ell}z_{\ell}\right)$
7:until $\hat{\theta}$ converges
8:
9:Output: Posterior mean $\hat{\theta}$ and covariance $\hat{\Sigma}$
Generalized linear models (GLMs) [36] extend linear models (Section 2.1) to non-linear monotone mean functions $g$ in (1). For instance, in logistic regression, $g(u)=1/(1+\exp[-u])$ is a sigmoid. The likelihood of observations in GLMs has the following form [25].
@@
Let $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ be a history of $N$ observations under mean function $g$ and the corresponding noise. Then $\log p(h\mid\theta_{*})\propto\sum_{\ell=1}^{N}y_{\ell}\phi_{\ell}^{\top}\theta_{*}-b(\phi_{\ell}^{\top}\theta_{*})+c(y_{\ell})$, where $c$ is a real function, and $b$ is a function whose derivative is the mean function, $\dot{b}=g$.
The posterior distribution of $\theta_{*}$ in GLMs does not have a closed form in general [10]. Therefore, it is often approximated by the Laplace approximation. Let the prior distribution of the model parameter be $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$, as in Section 2.1. Then the Laplace approximation is $\mathcal{N}(\hat{\theta},\hat{\Sigma})$, where $\hat{\theta}$ is the maximum a posteriori (MAP) estimate of $\theta_{*}$ and $\hat{\Sigma}$ is the corresponding covariance.
The MAP estimate $\hat{\theta}$ can be obtained by iteratively reweighted least squares (IRLS) [51], which we present in Algorithm 1. $\tt IRLS$ is an iterative method where the MAP estimate $\hat{\theta}$ is updated (line 5) until convergence. This is guaranteed under mild conditions due to the convexity of the problem. In the updates, $\dot{g}$ is the derivative of the mean function and the pseudo-observation $z_{\ell}$ (line 4) plays the role of the observation $y_{\ell}$ in the empirical mean $\bar{\theta}$. The IRLS solution has a similar structure to (3). Specifically, $\mathcal{N}(\hat{\theta},\hat{\Sigma})$ is a product of two multivariate Gaussians, representing the prior knowledge about $\theta_{*}$ and empirical evidence.
@@ 2.3 Towards Diffusion Model Priors
The assumption that $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$ is limiting, for instance because it precludes multimodal priors. We relax it by representing $p(\theta_{*})$ by a diffusion model, which we call a diffusion model prior. We propose an efficient posterior sampling approximation for this prior where the prior and empirical evidence are mixed similarly to (3) and $\tt IRLS$ (Section 4). Before we do this, we review diffusion models in Section 3.
@@ 3 Diffusion Models
A diffusion model [45, 18] is a generative model trained by diffusing samples from an unknown distribution. While multiple types of diffusion models exist [48], we adopt the probabilistic formulation and presentation of Ho et al. [18]. A diffusion model is a graphical model with $T$ stages indexed by $t\in[T]$. Each stage $t$ is associated with a latent variable $S_{t}\in\mathbb{R}^{d}$. A sample from the model is represented by an observed variable $S_{0}\in\mathbb{R}^{d}$. We visualize the diffusion model in Figure 1. In the forward process, the sample $s_{0}$ is diffused through a sequence of variables $S_{1},\dots,S_{T}$. This process is used to learn the reverse process, where the sample $s_{0}$ is generated through a sequence of variables $S_{T},\dots,S_{0}$. To sample $s_{0}$ from the posterior (Section 4), we add a random variable $H$, which represents partial information about $s_{0}$. We introduce forward and reverse diffusion processes next. Learning of the reverse process is described in Appendix B. While this is a critical component of diffusion models, it is not necessary to introduce our posterior approximations.
Forward process. The forward process diffuses an observed sample $s_{0}$ through a chain of latent variables $S_{1},\dots S_{T}$ (Figure 1). We denote the probability measure under this process by $q$ and define its joint probability distribution as
@@
where $q(s_{t}\mid s_{t-1})$ is the conditional density of mapping a less diffused $s_{t-1}$ to a more diffused $s_{t}$. The diffusion rate is set by parameters $\alpha_{t}\in(0,1)$ and $\beta_{t}=1-\alpha_{t}$. We also define $\bar{\alpha}_{t}=\prod_{\ell=1}^{t}\alpha_{\ell}$. The forward process is sampled as follows. First, a clean sample $s_{0}$ is chosen. Then $S_{t}\sim q(\cdot\mid s_{t-1})$ are sampled, starting from $t=1$ to $t=T$.
Reverse process. The reverse process maps a diffused sample $S_{T}$ to the observed sample $S_{0}$ through a chain of variables $S_{T},\dots,S_{0}$ (Figure 1). We denote the probability measure under this process by $p$ and define its joint probability distribution as
@@
where $p(s_{t-1}\mid s_{t})$ represents the conditional density of mapping a more diffused sample $s_{t}$ to a less diffused sample $s_{t-1}$. The function $\mu_{t}$ predicts the mean of $S_{t-1}\mid s_{t}$ and is learned from the data (Appendix B). As in Ho et al. [18], we keep the covariance fixed at $\Sigma_{t}=\tilde{\beta}_{t}I_{d}$, where $\tilde{\beta}_{t}$ is given in (13) in Appendix B. This is also called stable diffusion. We make this assumption only to simplify exposition. All our derivations in Section 4 hold when $\Sigma_{t}$ is learned, for instance as in Bao et al. [8].
The process is called reverse because it is learned by reversing the forward process. The reverse process is sampled as follows. First, a diffused sample $S_{T}\sim p$ is chosen. After that, $S_{t-1}\sim p(\cdot\mid s_{t})$ are sampled, starting from $t=T$ to $t=1$.
@@
$S_{T}\leftarrow S_{T-1}\leftarrow\dots\leftarrow S_{1}\leftarrow S_{0}\to H$ $S_{T}\to S_{T-1}\to\dots\to S_{1}\to S_{0}\to H$

### v2

@@ Abstract
Posterior sampling in contextual bandits with a Gaussian prior can be implemented exactly or approximately using the Laplace approximation. The Gaussian prior is computationally efficient but it cannot describe complex distributions. In this work, we propose approximate posterior sampling algorithms for contextual bandits with a diffusion model prior. The key idea is to sample from a chain of approximate conditional posteriors, one for each stage of the reverse diffusion process, which are obtained by the Laplace approximation. Our approximations are motivated by posterior sampling with a Gaussian prior, and inherit its simplicity and efficiency. They are asymptotically consistent and perform well empirically on a variety of contextual bandit problems.
@@ 1 Introduction
A multi-armed bandit [27, 6, 30] is an online learning problem where an agent sequentially interacts with an environment over $n$ rounds with the goal of maximizing its rewards. In each round, it takes an action and receives its stochastic reward. The mean rewards of the actions are unknown a priori and must be learned. This leads to the exploration-exploitation dilemma: explore actions to learn about them or exploit the action with the highest estimated reward. Bandits have been successfully applied to problems where uncertainty modeling and adaptation are beneficial, such recommender systems [32, 54, 25, 35] and hyper-parameter optimization [34].
@@
Contextual bandits [29, 32] with linear [13, 1] and generalized linear models (GLMs) [17, 33, 2, 26] have become popular due to the their flexibility and efficiency. The features in these models can be hand-crafted or learned from historic data [40], and the models can be also updated incrementally [1, 24]. While the original algorithms for linear and GLM bandits were based on upper confidence bounds (UCBs) [13, 1, 17], Thompson sampling (TS) is more popular in practice [11, 3, 43, 42]. The key idea in TS is to explore by sampling from the posterior distribution of model parameter $\theta_{*}$. TS uses the prior knowledge about $\theta_{*}$ to speed up exploration [11, 40, 36, 9, 22, 20, 4]. When the prior is a multivariate Gaussian, the posterior of $\theta_{*}$ can be updated and sampled from efficiently [11]. This prior has a limited expressive power, because it cannot even represent multimodal distributions. To address this, we study posterior sampling with a diffusion prior. The main benefit of such priors is that they can represent complex distributions and be learned from data.
@@
We make the following contributions. First, we propose novel posterior sampling approximations for linear models and GLMs with a diffusion model prior. The key idea is to sample from a chain of approximate conditional posteriors, one for each stage of the reverse process, which are estimated in a closed form. In linear models, each conditional is a product of two Gaussians, representing prior knowledge and diffused evidence (Theorem˜2). In GLMs, each conditional is obtained by a Laplace approximation, which mixes prior knowledge and evidence (Theorem˜4). Our approximations are motivated by posterior sampling with Gaussian priors, and inherit its simplicity and efficiency. Prior works (Section˜7) sampled from the posterior using the likelihood score, and their approximations become unstable when the score is high. We combine the likelihood with conditional priors, in each stage of the diffusion model, using the Laplace approximation. The resulting posterior concentrates at a single point and can be sampled from efficiently even if the likelihood score is high. We prove that this approximation is asymptotically consistent.
@@
Our second contribution is in theory. We properly derive our posterior approximations (Theorems˜2 and 4) and show their asymptotic consistency (Theorem˜3). The key idea in the proof of Theorem˜3 is that the conditional posteriors concentrate at a scaled unknown model parameter as the number of observations increases. While this claim is asymptotic, it is an expected property of a posterior distribution. Many prior works, such as Chung et al. [12], do not propose asymptotically consistent approximations. All of our main results rely on a novel approximation of clean samples by scaled diffused samples (Section˜4.3). The most challenging part of the analysis is Theorem˜3, where we analyze an asymptotic behavior of a chain of $T$ dependent random vectors.
@@
Our last contribution is an empirical evaluation on contextual bandits. We focus on bandits because the ability to represent all levels of uncertainty precisely is critical for exploration. Our experiments show that a score-based method fails to do so (Section˜6.2). Note that our posterior approximations are general and not restricted to bandits.
@@ 2 Setting
We start with introducing our notation. Random variables are capitalized, except for Greek letters like $\theta$. We denote the marginal and conditional probabilities under probability measure $p$ by $p(X=x)$ and $p(X=x\mid Y=y)$, respectively. When the random variables are clear from context, we write $p(x)$ and $p(x\mid y)$. We denote by $X_{n:m}$ and $x_{n:m}$ a collection of random variables and their values, respectively. For a positive integer $n$, we define $n]=\left\{1,\dots,n\right\}$. The indicator function is $\mathds{1}\!\left\{\cdot\right\}$. The $i$-th entry of vector $v$ is $v_{i}$. If the vector is already indexed, such as $v_{j}$, we write $v_{j,i}$. We denote the maximum and minimum eigenvalues of matrix $M\in\mathbb{R}^{d\times d}$ by $\lambda_{1}(M)$ and $\lambda_{d}(M)$, respectively.
@@
where $g:\mathbb{R}\to\mathbb{R}$ is the mean function and $\varepsilon_{\ell}$ is an independent zero-mean $\sigma^{2}$-sub-Gaussian noise for $\sigma>0$. Let $p(h\mid\theta_{*})$ be the likelihood of observations in history $h$ under model parameter $\theta_{*}$ and $p(\theta_{*})$ be its prior probability. By Bayes’ rule, the posterior distribution of $\theta_{*}$ given $h$ is
@@ 2.1 Linear Model
The posterior of $\theta_{*}$ in linear models can be derived as follows.
@@
Let $g$ in (1) be an identity and $\varepsilon_{\ell}\sim\mathcal{N}(0,\sigma^{2})$. Then the likelihood of $h$ under model parameter $\theta_{*}$ is $p(h\mid\theta_{*})\propto\exp[-\sum_{\ell=1}^{N}(y_{\ell}-\phi_{\ell}^{\top}\theta_{*})^{2}/(2\sigma^{2})]$.
[…]
$\textstyle\hat{\theta}=\hat{\Sigma}\left(\Sigma_{0}^{-1}\theta_{0}+\sigma^{-2}\sum_{\ell=1}^{N}\phi_{\ell}y_{\ell}\right)\,,\quad\hat{\Sigma}=\left(\Sigma_{0}^{-1}+\sigma^{-2}\sum_{\ell=1}^{N}\phi_{\ell}\phi_{\ell}^{\top}\right)^{-1}\,$,
are the posterior mean and covariance, respectively. In this work, we write them equivalently as
@@
where $\bar{\theta}=\sigma^{-2}\bar{\Sigma}\sum_{\ell=1}^{N}\phi_{\ell}y_{\ell}$ and $\bar{\Sigma}^{-1}=\sigma^{-2}\sum_{\ell=1}^{N}\phi_{\ell}\phi_{\ell}^{\top}$ are the empirical mean and inverse of its covariance, respectively. Therefore, the posterior of $\theta_{*}$ is a product of two multivariate Gaussians: $\mathcal{N}(\theta_{0},\Sigma_{0})$ representing prior knowledge about $\theta_{*}$ and $\mathcal{N}(\bar{\theta},\bar{\Sigma})$ representing empirical evidence.
@@ 2.2 Generalized Linear Model
Algorithm 1 $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt IRLS$: Iteratively reweighted least squares.
@@
5:  for stage $\ell=1,\dots,N$ do
[…]
7:  $\hat{\Sigma}\leftarrow\left(\Sigma_{0}^{-1}+\sum_{\ell=1}^{N}\dot{g}(\phi_{\ell}^{\top}\hat{\theta})\phi_{\ell}\phi_{\ell}^{\top}\right)^{-1}$
8:  $\hat{\theta}\leftarrow\hat{\Sigma}\left(\Sigma_{0}^{-1}\theta_{0}+\sum_{\ell=1}^{N}\dot{g}(\phi_{\ell}^{\top}\hat{\theta})\phi_{\ell}z_{\ell}\right)$
9:until $\hat{\theta}$ converges
10:
11:Output: Posterior mean $\hat{\theta}$ and covariance $\hat{\Sigma}$
Generalized linear models (GLMs) [37] extend linear models (Section˜2.1) to non-linear monotone mean functions $g$ in (1). For instance, in logistic regression, $g(u)=1/(1+\exp[-u])$ is a sigmoid. The likelihood of observations in GLMs has the following form [26].
@@
Let $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ be a history of $N$ observations under mean function $g$ and the corresponding noise. Then $\log p(h\mid\theta_{*})\propto\sum_{\ell=1}^{N}y_{\ell}\phi_{\ell}^{\top}\theta_{*}-b(\phi_{\ell}^{\top}\theta_{*})+c(y_{\ell})$, where $c$ is a real function and $b$ is a function whose derivative is the mean function, $\dot{b}=g$.
[…]
The MAP estimate $\hat{\theta}$ can be obtained by iteratively reweighted least squares (IRLS) [52], which we present in Algorithm˜1. $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt IRLS$ is a Newton-type algorithm that computes $\hat{\theta}$ iteratively (lines 6 and 7). It converges to the optimal solution due to the strong convexity of the problem. The solution has a similar structure to (3). That $\mathcal{N}(\hat{\theta},\hat{\Sigma}) is$ a product of two multivariate Gaussians, representing prior knowledge about $\theta_{*}$ and empirical evidence. The new quantities in GLMs are the derivative of the mean function $\dot{g}$ and pseudo-observations $z_{\ell}$ (line 5), which play the role of observations $y_{\ell}$ in Section˜2.1.
@@ 2.3 Towards Diffusion Model Priors
The assumption that $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$ is limiting, for instance because it precludes multimodal priors. We relax it by representing $p(\theta_{*})$ by a diffusion model, which we call a diffusion model prior. We propose efficient posterior sampling approximations for this prior, where the prior and empirical evidence are mixed similarly to (3) and $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt IRLS$. We review diffusion models next.
@@ 3 Diffusion Models
Diffusion models [46, 19] are generative models trained by diffusing samples from unknown and hard to represent distributions. They can be viewed in multiple ways [49]. We adopt the probabilistic formulation and presentation of Ho et al. [19]. A diffusion model is a graphical model with $T$ stages indexed by $t\in[T]$. Each stage $t$ is associated with a latent variable $S_{t}\in\mathbb{R}^{d}$. A sample from the model is represented by an observed variable $S_{0}\in\mathbb{R}^{d}$. We visualize a diffusion model in Figure˜1. In the forward process, a clean sample $s_{0}$ is diffused through a sequence of variables $S_{1},\dots,S_{T}$. This process is used to learn the reverse process, where the clean sample $s_{0}$ is generated through a sequence of variables $S_{T},\dots,S_{0}$. To sample $s_{0}$ from the posterior (Section˜4), we add a random variable $H$ that represents partial information about $s_{0}$. We introduce forward and reverse diffusion processes next. Learning of the reverse process is described in Appendix˜B. While this is a critical component of diffusion models, it is not necessary to introduce our posterior approximations.
Forward process. In the forward process, a clean sample $s_{0}$ is diffused through a chain of latent variables $S_{1},\dots S_{T}$ (Figure˜1). We denote the probability measure under this process by $q$ and define its joint probability distribution as
@@
where $q(s_{t}\mid s_{t-1})$ is the conditional density of mapping a less diffused $s_{t-1}$ to a more diffused $s_{t}$. The diffusion rate is set by parameters $\alpha_{t}\in(0,1)$ and $\beta_{t}=1-\alpha_{t}$. The forward process is sampled from as follows. First, a clean sample $s_{0}$ is chosen. Then $S_{t}\sim q(\cdot\mid s_{t-1})$ are sampled, from $t=1$ to $t=T$.
Reverse process. In the reverse process, a clean sample $s_{0}$ is generated through a chain of variables $S_{T},\dots,S_{0}$ (Figure˜1). We denote the probability measure under this process by $p$ and define its joint probability distribution as
@@
where $p(s_{t-1}\mid s_{t})$ is the conditional density of mapping a more diffused $s_{t}$ to a less diffused $s_{t-1}$. The function $\mu_{t}$ predicts the mean of $S_{t-1}\mid s_{t}$ and is learned (Appendix˜B). As in Ho et al. [19], we keep the covariance fixed at $\Sigma_{t}=\tilde{\beta}_{t}I_{d}$, where $\tilde{\beta}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{\ell=1}^{t}\alpha_{\ell}$. This is known as a stable diffusion. We make this assumption only to simplify exposition. All our derivations in Section˜4 hold when $\Sigma_{t}$ is learned, for instance as in Bao et al. [8].
This process is called reverse because it is learned by reversing the forward process. The reverse process is sampled from as follows. First, an initial diffused sample $S_{T}\sim p$ is sampled. After that, $S_{t-1}\sim p(\cdot\mid s_{t})$ are sampled, from $t=T$ to $t=1$.
@@
$S_{T}\leftarrow S_{T-1}\leftarrow\dots\leftarrow S_{1}\leftarrow S_{0}$ $S_{T}\to S_{T-1}\to\dots\to S_{1}\to S_{0}\to H$